how much water do you have to mix with 6 pints of 70% solution to produce a solution that is 23%?
Answer by stanbon(75887)

how much water do you have to mix with 6 pints of 70% solution to produce a solution that is 23%?
------
Equation:
water + water = water
0.30*6 + 1*x = 0.77(6+x)
-------
30*6 + 100x = 77*6 + 77x
23x = 47*6
x = 12.26 pints (amt. of water to add)
=========================================
Cheers,
Stan H.